Not logged in
Tue Jul 1 12:51:34 2008
Rana, Rob, Yoichi, John
The recent computer problems and MZ work had disturbed the alignment of the interferometer.
We adjusted the MC alignment back to nominal positions using old OSEM values. We then walked
the input beam to match the MC. Coupling into the interferometer has increased noticeably.
The rest of the IFO was then aligned to the new input beam.
Proceeding to full IFO locking we were able to engage the AO path and hand off CARM to SPOBDC.
Arm powers got up to 4.
If the new alignment proves successful we will centre all QPDs etc so we can easily return to
Uploaded Tue Jul 1 14:06:08 2008